# Croppable Knowledge Graph Embedding

### v1

@@ Abstract
Knowledge Graph Embedding (KGE) is a common method for Knowledge Graphs (KGs) to serve various artificial intelligence tasks. The suitable dimensions of the embeddings depend on the storage and computing conditions of the specific application scenarios. Once a new dimension is required, a new KGE model needs to be trained from scratch, which greatly increases the training cost and limits the efficiency and flexibility of KGE in serving various scenarios. In this work, we propose a novel KGE training framework MED, through which we could train once to get a croppable KGE model applicable to multiple scenarios with different dimensional requirements, sub-models of the required dimensions can be cropped out of it and used directly without any additional training. In MED, we propose a mutual learning mechanism to improve the low-dimensional sub-models performance and make the high-dimensional sub-models retain the capacity that low-dimensional sub-models have, an evolutionary improvement mechanism to promote the high-dimensional sub-models to master the knowledge that the low-dimensional sub-models can not learn, and a dynamic loss weight to balance the multiple losses adaptively. Experiments on 3 KGE models over 4 standard KG completion datasets, 3 real application scenarios over a real-world large-scale KG, and the experiments of extending MED to the language model BERT show the effectiveness, high efficiency, and flexible extensibility of MED.
@@ 1 Introduction
Knowledge Graphs (KGs) are composed of triples representing facts in the form of (head entity, relation, tail entity), abbreviated as (h, r, t). KG has been widely used in recommendation systems DBLP:conf/mm/ZhuZZYCZC21 ; DBLP:conf/icde/ZhangWYWZC21 , information extraction DBLP:conf/acl/HoffmannZLZW11 ; DBLP:conf/i-semantics/DaiberJHM13 , question answering DBLP:journals/corr/ZhangLHJLW016 ; DBLP:conf/www/DiefenbachSM18 and other tasks. A common way to apply a knowledge graph is to represent the entities and relations in the knowledge graph into continuous vector spaces, called knowledge graph embedding (KGE) DBLP:conf/nips/BordesUGWY13 ; DBLP:conf/iclr/SunDNT19 , and then use the vector representation of entities and relations to serve a variety of tasks.
KGEs with higher dimensions have greater expressive power and usually achieve better performance, but this also means a larger number of parameters and requires more storage space and computing resources DBLP:conf/wsdm/ZhuZCCC0C22 ; DBLP:conf/acl/Sachan20 . The appropriate dimensions of the KGE are different for different devices or scenarios. As shown in Fig. 1, large remote servers have large storage space and sufficient computing resources to support high-dimensional KGE with good performance, while small and medium-sized terminal devices, such as vehicle-mounted systems or smartphones, can only accept low-dimensional KGE due to limited computing power and storage capacity. Therefore, according to the conditions of different devices or scenes, people tend to train the KGE with appropriate dimensions and as high quality as possible. However, the challenge is that once a new dimension is required, a new KGE needs to be trained from scratch. Especially when only low-dimensional KGE can be applied, to ensure good performance, the additional model compression technology such as knowledge distillation DBLP:journals/corr/HintonVD15 ; DBLP:conf/wsdm/ZhuZCCC0C22 is needed during training. This significantly increases training costs and limits KGE’s efficiency and flexibility in serving different scenarios.
@@
Thus a new concept "croppable KGE" is proposed and we are interested in the research question that is it possible to train a croppable KGE, with which KGEs of various required dimensions can be cropped out of it, directly be used without any additional training, and achieve promising performance?
@@
In this work, our main idea of croppable KGE learning is to train an entire KGE that contains many sub-models of different dimensions in it. These sub-models share their embedding parameters and are trained simultaneously. The goal is that the low-dimensional sub-models can benefit from the more expressive high-dimensional sub-models, while the high-dimensional sub-models retain the ability of the low-dimensional sub-models and master the knowledge that the low-dimensional sub-models cannot. Based on this idea, we propose a croppable KGE training framework MED, which consists of three main modules, the Mutual learning mechanism, the Evolutionary improvement mechanism, and the Dynamic loss weight to achieve the above purpose. Specifically, the mutual learning mechanism is based on knowledge distillation and it makes pairwise neighbor sub-models learn from each other, so that the performance of the lower-dimensional sub-model can be improved, and the higher-dimensional sub-model can retain the ability of the lower-dimensional sub-model. The evolutionary improvement mechanism helps the high-dimensional sub-model master more knowledge that the low-dimensional sub-model cannot by making the high-dimensional sub-model pay more attention to learn the triples that the low-dimensional sub-model can’t correctly predict. The dynamic loss weight is designed to adaptively balance multiple losses of different sub-models according to their dimensions and further improve the overall performance.
@@
We evaluate the effectiveness of our proposed MED by implementing it on three typical KGE methods and four standard KG datasets. We also prove its practical value by applying MED to a real-world large-scale KG and downstream tasks. Furthermore, we demonstrate the extensibility of MED by implementing it on language model BERT DBLP:conf/naacl/DevlinCLT19 and GLUE DBLP:conf/iclr/WangSMHLB19 benchmarks. The experimental results show that (1) MED successfully trains a croppable KGE model available for various dimensional requirements, which contains multiple parameter-shared sub-models of different dimensions that of high performance and can be used directly without additional training; (2) the training efficiency of MED is far higher than that of independently training multiple KGE models of different sizes or obtaining them by knowledge distillation. (3) MED can be flexibly extended to other neural network models besides KGE and achieve good performance; (4) our proposed mutual learning mechanism, evolutionary improvement mechanism, and dynamic loss weight are effective and necessary for MED to achieve overall optimal performance. In summary, our contributions are as follows:
  • We propose a new research question and task: training croppable KGE, from which KGEs of different dimensions can be cropped and used directly without any additional training.
@@
  • We propose a novel framework MED, including a mutual learning mechanism, an evolutionary improvement mechanism, and a dynamic loss weight, to ensure the overall performance of all sub-models during training the croppable KGE.
@@
  • We experimentally prove that all sub-models of MED work well, especially the performance of the low-dimensional sub-models exceeding the KGE with the same dimension trained by the state-of-the-art distillation-based methods. MED also shows excellent performance in real-world applications and good extensibility on other types of neural networks.
@@ 2 Related Work
This work is to achieve a croppable KGE that meets different dimensional requirements. One of the most common methods to obtain a good-performance KGE of the target dimension is utilizing knowledge distillation with a high-dimensional powerful teacher KGE. Thus, we focus on two research fields most relevant to our work: knowledge graph embedding and knowledge distillation.
@@ 2.1 Knowledge Graph Embedding
Knowledge graph embedding (KGE) technology has been widely applied with the key idea of mapping entities and relations of a KG into continuous vector spaces as vector representations, which can further serve various KG downstream tasks. TransE DBLP:conf/nips/BordesUGWY13 is the most representative translation-based KGE method by regarding the relation as a translation from the head to tail entity. Variants of TransE include TransH DBLP:conf/aaai/WangZFC14 , TransR DBLP:conf/aaai/LinLSLZ15 , TransD DBLP:conf/acl/JiHXL015 and so on. RESCAL DBLP:conf/icml/NickelTK11 is the first one based on vector decomposition, and then to improve it, DistMult DBLP:journals/corr/YangYHGD14a , ComplEx DBLP:conf/icml/TrouillonWRGB16 , and SimplE DBLP:conf/nips/Kazemi018 are proposed. RotatE DBLP:conf/iclr/SunDNT19 is a typical rotation-based method that regards the relation as the rotation between the head and tail entities. QuatE DBLP:conf/nips/0007TYL19 and DihEdral DBLP:conf/acl/XuL19 work with a similar idea. PairRE DBLP:conf/acl/ChaoHWC20 uses two relation vectors to project the head and tail entities into an Euclidean space to encode complex relational patterns. With the development of neural networks, KGEs based on graph neural networks (GNNs) DBLP:conf/aaai/DettmersMS018 ; DBLP:conf/naacl/NguyenNNP18 ; DBLP:conf/esws/SchlichtkrullKB18 ; DBLP:conf/iclr/VashishthSNT20 are also proposed. Although the KGEs are simple and effective, there is an obvious challenge: In different scenarios, the required KGE dimensions are different, which depends on the storage and computing resources of the device. It has to train a new KGE model from scratch for a new dimension requirement, which greatly increases the training cost and limits the flexibility for KGE to serve diversified scenarios.
@@ 2.2 Knowledge Distillation
High-dimensional KGEs have strong expression ability due to the large number of parameters, but require a lot of storage and computing resources, and are not suitable for all scenarios, especially small devices. To solve this problem, a common way is to compress a high-dimensional KGE to the target low-dimensional KGE by knowledge distillation DBLP:journals/corr/HintonVD15 ; DBLP:conf/aaai/MirzadehFLLMG20 and quantization DBLP:conf/acl/BaiZHSJJLLK20 ; DBLP:conf/iclr/StockFGGGJJ21 technology.
Quantization replaces continuous vector representations with lower-dimensional discrete codes. TS-CL DBLP:conf/acl/Sachan20 is the first work of KGE compression applying quantization. LightKG DBLP:conf/cikm/WangWLG21 uses a residual module to induce diversity among codebooks. However, quantization cannot improve the inference speed so it’s still not suitable for devices with limited computing resources.
Knowledge distillation (KD) has been widely used in Computer Vision DBLP:conf/aaai/MirzadehFLLMG20 and Natural Language Processing DBLP:conf/naacl/DevlinCLT19 ; DBLP:conf/emnlp/SunCGL19 , helping reduce the model size and increase the inference speed. The core idea is to use the output of a large teacher model to guide the training of a small student model. DualDE DBLP:conf/wsdm/ZhuZCCC0C22 is a representative KD-based work to transfer the knowledge of high-dimensional KGE to low-dimensional KGE. It considers the mutual influences between the teacher and student and finetunes the teacher during training.MulDE DBLP:conf/www/Wang0MS21 transfers the knowledge from multiple low-dimensional teacher models to a student model for hyperbolic KGE. ISD DBLP:journals/corr/abs-2206-02963 improves low-dimensional KGE by making it play the teacher and student roles alternatively during training. Among these methods, DualDE DBLP:conf/wsdm/ZhuZCCC0C22 is more relevant to our work, both have the setting of high-dimensional teacher and low-dimensional student models. In this work, we propose a novel KD-based KGE training framework MED, one training can obtain a croppable KGE that meets multiple dimensional requirements.
@@ 3 Preliminary
KGE method Scoring Function $f(\mathbf{h},\mathbf{r},\mathbf{t})$
TransE DBLP:conf/nips/BordesUGWY13 $-\left\|\mathbf{h}+\mathbf{r}-\mathbf{t}\right\|$
RotatE DBLP:conf/iclr/SunDNT19 $-\left\|\mathbf{h}\circ\mathbf{r}-\mathbf{t}\right\|$
PairRE DBLP:conf/acl/ChaoHWC20 $-\left\|\mathbf{h}\circ\mathbf{r}^{H}-\mathbf{t}\circ\mathbf{r}^{T}\right\|$
Knowledge graph embedding (KGE) methods aim to express the relations between entities in a continuous vector space through a scoring function $f$. Specifically, given a knowledge graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$ where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ are the sets of entities, relations and all observed triples, we utilize the triple scoring function to measure the plausibility of triples in the embedding space for a triple $(h,r,t)$ where $h\in\mathcal{E},r\in\mathcal{R}$ and $t\in\mathcal{E}$. The triple score function is denoted as $s_{(h,r,t)}=f(\mathbf{h},\mathbf{r},\mathbf{t})$ with embeddings of head entity h, relation r and tail entity t as input. Table 1 summarizes the scoring functions of some popular KGE methods, where $\circ$ is the Hadamard product. The higher the triple score, the more likely the model is to judge the triples as true.
@@ 4 MED Framework
As shown in Fig. 2, our croppable KGE framework MED contains multiple (let’s say $n$) sub-models of different dimensions in it, denoted as $M_{i}(i=1,2...,n)$ with dimension of $d_{i}$. Each sub-model
@@
MED is based on knowledge distillation DBLP:journals/corr/HintonVD15 ; DBLP:journals/corr/abs-1903-12136 ; DBLP:conf/naacl/DevlinCLT19 technique that the student learns by fitting the hard (ground-truth) label and the soft label from the teacher simultaneously. In MED, we first propose a mutual learning mechanism that makes low-dimensional sub-models learn from high-dimensional sub-models to achieve better performance, and makes high-dimensional sub-models also learn from low-dimensional sub-models to retain the abilities that low-dimensional sub-models already have. Then, we propose an evolutionary improvement mechanism to enable high-dimensional sub-models to master the knowledge that the low-dimensional sub-models can not learn well. Finally, we train MED with dynamic loss weight to adaptively balance multiple optimization objectives of sub-models.
@@ 4.1 Mutual Learning Mechanism
We treat each sub-model $M_{i}$ as the student of its higher-dimensional neighbor sub-model $M_{i+1}$ to achieve better performance, since high-dimensional KGEs usually have more expressive power than low-dimensional ones due to more parameters DBLP:conf/acl/Sachan20 ; DBLP:conf/wsdm/ZhuZCCC0C22 . We also treat sub-model $M_{i}$ as the student of its lower-dimensional neighbor sub-model $M_{i-1}$, so the higher-dimensional sub-model can review what the lower-dimensional sub-model has learned and retain the low-dimensional one’s existing abilities. Thus, pairwise neighbor sub-models serve as both teachers and students, learning from each other. The mutual learning loss between each pair of neighbor sub-models is
@@
where $s_{(h,r,t)}^{i}$ is the score of triple $(h,r,t)$ output by sub-model $M_{i}$ and reflects the possibility that this triplet exists, $\mathcal{T}^{-}=\mathcal{E}\times\mathcal{R}\times\mathcal{E}\setminus\mathcal{T}$ is the negative triple set, $n$ is the number of sub-models, and $d_{\delta}$ is Huber loss Huber1964Robust with $\delta=1$ commonly used in knowledge distillation for KGE  DBLP:conf/wsdm/ZhuZCCC0C22 . MED makes each sub-model only learn from its neighbor sub-models. The advantage is that this not only reduces the computational complexity of training but also makes every pair of teacher and student models have a relatively small dimension gap, which is important and effective because the large gap of dimensions between teacher and student will destroy the distillation effect DBLP:conf/aaai/MirzadehFLLMG20 ; DBLP:conf/wsdm/ZhuZCCC0C22 .

### v2

@@ Abstract
Knowledge Graph Embedding (KGE) is a common approach for Knowledge Graphs (KGs) in AI tasks. Embedding dimensions depend on application scenarios. Requiring a new dimension means training a new KGE model from scratch, increasing cost and limiting efficiency and flexibility. In this work, we propose a novel KGE training framework MED. It allows one training to obtain a croppable KGE model for multiple scenarios with different dimensional needs. Sub-models of required dimensions can be directly cropped and used without extra training. In MED, we propose a mutual learning mechanism to improve the low-dimensional sub-models and make high-dimensional sub-models retain the low-dimensional sub-models’ capacity, an evolutionary improvement mechanism to promote the high-dimensional sub-models to master the triple that the low-dimensional sub-models can not, and a dynamic loss weight to adaptively balance the multiple losses. Experiments on 4 KGE models across 4 standard KG completion datasets, 3 real-world scenarios using a large-scale KG, and extending MED to the BERT language model demonstrate its effectiveness, high efficiency, and flexible extensibility.
@@ 1 Introduction
Knowledge Graphs (KGs) consist of triples in the form of (head entity, relation, tail entity), abbreviated as (h, r, t). KGs are widely used in recommendation systems Zhu et al. (2021), information extraction Daiber et al. (2013), question answering Diefenbach et al. (2018), etc. A common KG application way is knowledge graph embedding (KGE) Bordes et al. (2013); Sun et al. (2019b), which maps KG entities and relations into continuous vector spaces for various tasks.
@@
Higher-dimensional KGEs have more expressive power and better performance but come with more parameters, demanding more storage and computing resources Zhu et al. (2022); Sachan (2020). The suitable KGE dimensions vary by device or scenario. As Fig. 1 shows, large remote servers can support high-dimensional KGEs, while small-to-medium terminal devices like in-vehicle systems or smartphones can only handle low-dimensional ones due to resource limits. Thus, people prefer training high-quality KGEs with appropriate dimensions. However, a new KGE must be trained from scratch when a new dimension is needed. Especially for low-dimensional KGEs, model compression techniques like knowledge distillation Hinton et al. (2015); Zhu et al. (2022) are required to ensure performance. This raises training costs and restricts KGE’s efficiency and flexibility across scenarios.
@@
Thus a new concept "croppable KGE" is proposed and we are interested in the research question that is it possible to train a croppable KGE, with which KGEs of various required dimensions can be cropped out of it and directly used without extra training, and perform well?
@@
In this work, our croppable KGE learning idea is to train an entire KGE with multiple different-dimensional sub-models sharing embedding parameters and trained simultaneously. The aim is for low-dimensional sub-models to benefit from high-dimensional ones, and high-dimensional sub-models to retain low-dimensional ones’ ability and learn what they can’t. Based on this, we propose a croppable KGE training framework MED, with three main modules: Mutual learning mechanism, Evolutionary improvement mechanism, and Dynamic loss weight. Specifically, the mutual learning mechanism, based on knowledge distillation, enables neighbor sub-models to learn from each other, improving low-dimensional sub-model performance and helping high-dimensional ones retain low-dimensional abilities. The evolutionary improvement mechanism helps high-dimensional focus on and master triples that low-dimensional ones can’t correctly predict. The dynamic loss weight adaptively balances sub-models’ losses according to dimension to enhance overall performance.
@@
We evaluate the effectiveness of MED by implementing it on four typical KGE methods and four standard KG datasets. We prove its practical value via a real-world large-scale KG and downstream tasks, and demonstrate its extensibility on BERT Devlin et al. (2019) and GLUE Wang et al. (2019) benchmarks. Experimental results show that: (1) MED trains a croppable KGE for various dimensional needs, with high-performing, parameter-shared sub-models ready for direct use. (2) MED has far higher training efficiency than independent training or knowledge distillation. (3) MED can be extended to other neural networks and performs well. (4) The three modules in MED are crucial for optimal performance. In summary, our contributions are as follows:
  • We introduce the research question and task of training croppable KGE, allowing direct use of differently-dimensional KGEs.
@@
  • We propose a novel framework MED, including a mutual learning mechanism, an evolutionary improvement mechanism, and a dynamic loss weight, to ensure the overall performance of all sub-models.
@@
  • We show that MED’s sub-models perform well, with low-dimensional ones outperforming KGEs trained by SOTA distillation methods. MED also shows good performance and extensibility in real-world applications and other types of neural networks.
@@ 2 Related Work
This work is to achieve a croppable KGE for different dimensional needs. A common way to get a good-performance KGE of the target dimension is using knowledge distillation with a high-dimensional powerful teacher KGE. So, we focus on two relevant research areas: knowledge graph embedding and knowledge distillation.
@@ 2.1 Knowledge Graph Embedding
Knowledge graph embedding (KGE) maps KG entities and relations into continuous vector spaces for downstream tasks. TransE Bordes et al. (2013), a representative translation-based KGE, treats relations as translations between entities. Its variants include TransHWang et al. (2014), TransRLin et al. (2015), TransDJi et al. (2015), etc. RESCALNickel et al. (2011), based on vector decomposition, was followed by improvements like DistMultYang et al. (2015), ComplExTrouillon et al. (2016), and SimplEKazemi and Poole (2018). RotatESun et al. (2019b), a rotation-based method, views relations as rotations between entities, similar to QuatEZhang et al. (2019) and DihEdralXu and Li (2019). PairREChao et al. (2021) uses two relation vectors for complex pattern encoding. With neural network development, KGEs based on graph neural networks (GNNs) Dettmers et al. (2018); Nguyen et al. (2018); Schlichtkrull et al. (2018); Vashishth et al. (2020) also emerged. While KGEs are simple and effective, a key challenge remains: Different scenarios demand different KGE dimensions based on device resources. Training a new KGE model from scratch for each dimension requirement hikes training costs and restricts flexibility in serving diverse scenarios.
@@ 2.2 Knowledge Distillation
High-dimensional KGEs possess strong expressive power thanks to numerous parameters. However, they demand substantial storage and computing resources, rendering them unsuitable for all scenarios, particularly those involving small devices. To address this, a prevalent approach is to compress a high-dimensional KGE into target dimension via knowledge distillation Hinton et al. (2015); Mirzadeh et al. (2020) and quantization Bai et al. (2021); Stock et al. (2021) technology.
Quantization replaces continuous vector representations with lower-dimensional discrete codes. TS-CL Sachan (2020) was the first to apply quantization in KGE compression. LightKG Wang et al. (2021a) uses a residual module to create diversity in codebooks. Yet, as quantization doesn’t boost inference speed, it’s still not suitable for devices with limited computing resources.
Knowledge distillation (KD) is widely used in Computer Vision Mirzadeh et al. (2020) and Natural Language Processing Devlin et al. (2019); Sun et al. (2019a) to shrink model size and boost inference speed. Its core is using a large teacher model’s output to guide a small student model’s training. DualDE Zhu et al. (2022) is a representative KD-based work to transfer knowledge from high- to low-dimensional KGE, considering teacher-student mutual influences. MulDE Wang et al. (2021b) transfers knowledge from multiple hyperbolic KGE models to one student. ISD Zhou et al. (2022b) improves low-dimensional KGE by playing teacher and student roles alternatively. IterDE Liu et al. (2023) introduces iterative distillation for smooth knowledge transfer. Other KG-related distillation works include PMD Fan et al. (2024), which applies distillation to pre-trained language models for KG completion, IncDE Liu et al. (2024), using distillation between same-dimensional models at different times for incremental learning, and SKDE Xu et al. (2024), which proposes self-knowledge distillation to avoid a complex teacher model. Among these, DualDE Zhu et al. (2022) and IterDE Liu et al. (2023) are most relevant to our work as they compress high-dimensional teacher into low-dimensional student. In this work, we propose a novel KD-based KGE training framework MED, one training can obtain a croppable KGE that meets multiple dimensional requirements.
@@ 3 Preliminary
KGE method Scoring Function $f(\mathbf{h},\mathbf{r},\mathbf{t})$
TransE Bordes et al. (2013) $-\left\|\mathbf{h}+\mathbf{r}-\mathbf{t}\right\|$
SimplE Kazemi and Poole (2018) $\frac{1}{2}(<h^{H},r,t^{T}>+<t^{H},r^{-1},h^{T}>)$
RotatE Sun et al. (2019b) $-\left\|\mathbf{h}\circ\mathbf{r}-\mathbf{t}\right\|$
PairRE Chao et al. (2021) $-\left\|\mathbf{h}\circ\mathbf{r}^{H}-\mathbf{t}\circ\mathbf{r}^{T}\right\|$
@@
Knowledge graph embedding (KGE) methods use a scoring function $f$ to represent entities and relations in a continuous vector space. Given a KG $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$ where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ are sets of entities, relations and observed triples, for a triple $(h,r,t)$ ($h\in\mathcal{E},r\in\mathcal{R}$, $t\in\mathcal{E}$), we use the triple scoring function $s_{(h,r,t)}=f(\mathbf{h},\mathbf{r},\mathbf{t})$ (with entity and relation embeddings as input) to measure triple plausibility in the embedding space. Table 1 summarizes scoring functions of some popular KGE methods, where $\circ$ is the Hadamard product, $<x^{1},...,x^{k}>=\sum_{i}x^{1}_{i}...x_{i}^{k}$ is the generalized dot product. Higher triple scores mean the model is more likely to consider triples true. The optimization objective of the KGE model is
$L_{KGE}=-\sum_{(h,r,t)\in\mathcal{T}\cup\mathcal{T}^{-}}y\log\sigma(s_{(h,r,t)})$ (1)
$+(1-y)\log(1-\sigma(s_{(h,r,t)})),$
where $\mathcal{T}^{-}=\mathcal{E}\times\mathcal{R}\times\mathcal{E}\setminus\mathcal{T}$ is negative triple set, $\sigma$ is Sigmoid function, $y$ is ground-truth label of $(h,r,t)$, $y=1$ for positive and 00 for negative one.
@@ 4.1 Mutual Learning Mechanism
We consider each sub-model $M_{i}$ as the student of its higher-dimensional neighbor $M_{i+1}$ for better performance, as high-dimensional KGEs with more parameters are more expressive Sachan (2020); Zhu et al. (2022). We also make $M_{i}$ the student of its lower-dimensional neighbor $M_{i-1}$, enabling the higher-dimensional sub-model to review and retain the lower-dimensional one’s capabilities. So, pairwise neighbor sub-models are both teacher and student, learning from each other. Mutual learning loss between each pair of neighbor sub-models is
@@
where $s_{(h,r,t)}^{i}$ is the score of triple $(h,r,t)$ from sub-model $M_{i}$, indicating the triple’s existence likelihood. $d_{\delta}$ is Huber loss Huber and Peter (1964) with $\delta=1$, often used in KGE knowledge distillationZhu et al. (2022). In MED, each sub-model learns only from its neighbor sub-models. This not only cuts down training computational complexity but also keeps the dimension gap between each teacher-student pair relatively small. A large dimension gap between them can undermine the distillation effect Mirzadeh et al. (2020); Zhu et al. (2022), making our approach crucial and effective.